The length of each side of square A is increased by 100% to make square B. If the length of the side of square B is increased by 50% to make square C, by what percent is the area of Square C greater than the sum of the areas of squares A and B?
Answer by edjones(8007)

Let x be a side of square a.
2x is the length of a side of square b.
3x is the length of a side of square c.
A[a]=x^2
A[b]=(2x)^2=4x^2
A[c]=(3x)^2=9x^2
9x^2/(x^2+4x^2)
=9x^2/5x^2
=9/5
=1.8
So, the area of triangle C is 80% greater than the sum of the areas of triangles A and B.
.
Ed